ON A CERTAIN DAY, THE NUMBER OF BRITISH POUNDS (P) THAT COULD BE OBTAINED FOR (E) EUROS IS MODELED BY P=1.48598E. ALSO THE NUMBER OF EUROS (E) THAT COULD BE OBTAINED FOR (D) US DOLLARS IS MODELED BY E=1.22660D. HOW MANY U.S. DOLLARS COULD BE OBTAINED FOR 1000 BRITISH POUNDS.
Answer by tvandenberg(45)

P = 1.48598 (1.23660 D) = 1.83756 D
-->
D = P / 1.83756 = 1000 / 1.83756 = 544